OpenStudy (anonymous):

Find the geometric mean of each pair of numbers. If necessary, give the answer in simplest radical form. 5 and 12

OpenStudy (kirbykirby):

The geometric mean would be: \(\sqrt{5\cdot12}=\sqrt{60}=2\sqrt{15}\)

OpenStudy (whpalmer4):

In case you're wondering how the other poster went from \(\sqrt{60}\) to \(2\sqrt{15}\): \[\sqrt{60} = \sqrt{4*15} = \sqrt{4}*\sqrt{15} = 2*\sqrt{15} = 2\sqrt{15}\]